OpenStudy (anonymous):

For their CD collections, Alan, Tom, and Barbara have a total of 52 CD’s. If Barbara had 6 more, then she would have as many as both of the boys together. Tom has three more CD’s than Alan. How many CD’s does each of the three have?

OpenStudy (kropot72):

We can write three equations using the given information. Let A be the number of CDs that Alan has, T be Tom's number of CDs and B be Barbara's number of CDs. A + T + B = 52 ..........(1) B + 6 = A + T ............(2) T - 3 = A ...................(3) Do you follow, so far?

OpenStudy (anonymous):

Yes:)

OpenStudy (kropot72):

Do you have any ideas on how to solve these equations?

OpenStudy (anonymous):

Breaking it down like that helped but I am still very lost when it comes to actually solving it.

OpenStudy (kropot72):

One way to solve is to rearrange (2) as follows: B + 6 = A + T ............(2) Rearranging we get A + T - B = 6 ............(4) Now if we add equations (1) and (4) can you find the result?

OpenStudy (kropot72):

Add these: A + T + B = 52 ..........(1) A + T - B = 6 ............(4)

OpenStudy (anonymous):

So would A + T equal 46? Then would I just divide by two then take 3 from A and add it to T?

OpenStudy (kropot72):

Adding (1) and (4) gives: 2A + 2T = 58 ...........(5) Do you follow?

OpenStudy (anonymous):

Oh, okay. Yes I'm following.

OpenStudy (kropot72):

Equation (3) gives A in terms of T A = T - 3 Therefore we can substitute for A in equation (5) giving: 2(T - 3) + 2T = 58 ..........(6) Can you solve (6) to find the value of T?

OpenStudy (anonymous):

So T=16?

OpenStudy (kropot72):

Correct. Good work! Now substitute your value for T into (3) to find the value of A: T - 3 = A ...................(3)

OpenStudy (anonymous):

A= 13 T=16 B=23 Correct?

OpenStudy (kropot72):

Correct. Well done!

OpenStudy (anonymous):

Alright, thank you so much for your help! :)

OpenStudy (kropot72):

You're welcome :)